PREX Meeting Notes, Feb 8, 2010

Present: Bob Michaels, Ole Hansen, Mindy Jen, John LeRose, 
         Zafar Ahmed, Rupesh Silwal, Dustin McNulty, Kent Paschke,
         Luis Mercado, Sasha Glamazdin, Megan Friend, Rich Holmes

REMINDER:   Collaboration meeting planned for Feb 19-20.  
Therefore no meeting on Feb 15.

Bob Michaels

  - PREX target being made (Phil Adderley). Measuring foil thicknesses, etc.  
  - Septum is on pivot.  Technician manpower very limited !  Could slow us down.
  - Space being made for GEM DAQ (thanks Bogdan and Albert)
  - Plan to install GEM DAQ on L-HRS this week with Alex's help.
     - will do R-HRS GEM DAQ when detector stack goes back in
  - Work on CompSimple scheme for Compton
     - combines CMU-FADC crate with the ETROC
     - runs at 240 Hz helicity flip rate
     - coincidences simulated with pulsers
     - simple analysis framework (merger of existing codes)

Moller (Sasha)

  - Magnet and target installed on beam line
  - Survey Group finished all surveys (field, targets, etc).
  - Exercising target motion
  - Will cool down after getting cryo-lines
  - On track and looking good !

Zafar and Mindy

  - Comparing their Moller rates and asymmetries.  It's converging.
  - Zafar discussed problems getting controlled rates with PMT
    (idea to use randoms to test the DAQ, also useful for Compton)
  - Dustin will help with the PMT problem.
  - Mindy and Zafar have both volunteered for Compton work.

John LeRose

  - Septum magnet looks good.
  - Earlier reported field asymmetries were artefacts of probe positioning.
  - Now looks very symmetric, agrees with TOSCA.
      (with thanks to Javier Gomez, Jack Segal, Mindy Jen, and Paul Brindza)
  - John to produce transfer functions for Podd (analyzer) and hamc (MC).

Luis Mercado

  - Helicity multiplet analysis and noise studies - see HAPLOG entries
  - Dramatic suppression of 60 Hz noise.
  - Slightly wrong helicity pattern 
      - Kent thinks we could live with it, but Qweak will want it fixed.
      - Riad Suleiman working on fixing it.
  - Still see noise, probably the LED, will investigate
  (extensive conversation with Kent suggesting tests such as
   turning off LED, see h.c. diffs, cut on pair synch, etc.)
Dustin McNulty

   - SU scint checkout with Zafar :
     - Long cosmics runs for pulse-height spectra.  Looks good.
     - Will check again after 15m cables, fanout, and amplifier
     - Finished beam-right scint, next will do beam-left scint.
     - To install with Sasha's guidance.

   - New dithering :
     - Func. generator being arranged with ACC-div support.
     - Need cables to BSY.  Discussion of who/when.
     - Need to modify BMW software (Dustin, Kent)
     - question re: post-pan analysis ?  TBD.
   - HAMC simulation of PREX focal plane distrib.
     - for A_T detector and for thin targets
     - trying to finalize detector designs.

   - Wants to work with main detectors and GEMS in hall, soon.
     - preparing trigger electronics, and general cleanup
     - discussed relay switch: likely the old one is too noisy,
            may need accesses to switch counting to integrating
   - question: can we block A_T hole at high current ?
          (need a long arm ...)

Ole Hansen

   - GEM Analysis software making good progress.  
        Based on existing bigbite code.
        Decoder - done
        Weighting / averaging of hit positions
        Taking out L/R ambiguity resolution 
           (Bigbite remnant code that is not needed.)
        Need more coordinate directions (if I understood Ole)
        Need refurbished target reconstruction code.

Megan Friend

   - FADC pedestal shifts with laser state, see HAPLOG.
   - Crosstalk ?
   - Discussion of implications -- it wasn't clear.
   - Could check during downtime with various signals ?
   - CMU/FADC does not see the BCM-detector crosstalk (see R. Holmes below)
Kent Paschke

   - Getting summary of Optics data for HAPPEX-3 (Rupesh helping)
   - Soon (~1 week) need injector DAQ for beam studies (Joe Grames)
   - Laser table work (start ~2 weeks) 
      - Multi-element array for spot-size asy
      - Cables run to HAPPEX DAQ in ISB.
   - Discussed LED tests of new detector systems.
      - to have separate sub-group meeting on detectors and noise

Richard Holmes

   - See HAPLOG 2032 -- good summary of recent Compton issues.
   - Asymmetry systematic -- ostensibly crosstalk of detector to BCM
      - Might be a rate-dependent DAQ failure of unknown origin
      - Creates a false asymmetry in BCMs and side bumps.
      - Alex Camsonne to help investigate the hardware (no progress yet)

Note the Run Plan for the PREX run, to be discussed at the Collab. mtg Feb 19-20.  

Index of Meetings